# Reweight-annealing method for calculating the value of partition function via quantum Monte Carlo

### v2

@@ Abstract
On the other hand, the high-precision PF and its related quantities, such as free energy and entropy, are significantly important in numerical calculations since it means more effective information. For phase transitions no matter within or beyond the Ginzburg-Landau paradigm, an efficient and accurate way to straightforwardly obtain FP would be crucial, because its (high order) derivative reveals potential phase transitions. Besides, the precision of entropy can even reveal important intrinsic physics. In conformal field theory (CFT), the prefactor in front of the sub-leading term of entropy often carries universal information Calabrese and Lefevre (2008); Fradkin and Moore (2006); Nussinov and Ortiz (2009a, b); Casini and Huerta (2007); Ji and Wen (2019, 2020); Kong et al. (2020); Wu et al. (2021a); Zhao et al. (2021); Wu et al. (2021b). The thermal entropy also relates to the distinctiveness of many novel condensed matters, such as spin ice Snyder et al. (2001); Skjærvø et al. (2020); Udagawa et al. (2021); Bramwell and Harris (2020); Udagawa et al. (2021), topological order Jiang et al. (2012); Castelnovo et al. (2012); Sachdev (2018); Levin and Wen (2006); Yan et al. (2021, 2022), fracton phases Pretko et al. (2020); Nandkishore and Hermele (2019); Gromov and Radzihovsky (2024); Zhou et al. (2022), etc. Even a small numerical error may make our concerned quantities deviate a lot because the sub-leading term is much more sensitive than the leading one.
@@
In this paper, we present a more accurate and efficient method for calculating PF, dubbed reweight-annealing algorithm, within the framework of Monte Carlo (MC) simulation, either quantum or classical. It is inspired by thermal and quantum annealing algorithm in the quantum simulation area Huse and Fisher (1986); Santoro et al. (2002); Lucas (2014); Kadowaki and Nishimori (1998); Santoro et al. (2002); Mézard and Montanari (2009); Heim et al. (2015); Yan et al. (2023); Ding et al. (2023) and recent developments in the entanglement calculations of QMC Alba (2017); D’Emidio (2020); Zhao et al. (2022a, b); Pan et al. (2023); Zhou et al. (2024). Our method not only has high accuracy for calculating PF, but is unbiased, easy to implement and naturally can be parallelized on computer.
@@
Basic idea.- Given a partition function $Z(\beta)$, where $\beta$ is a general parameter which is not limited to temperature, but can be some other parameter in the Hamiltonian. For two parameter points $\beta^{\prime}$ and $\beta^{\prime\prime}$, we define the ratio
$\gamma(Z;\beta^{\prime},\beta^{\prime\prime})\equiv\frac{Z(\beta^{\prime\prime})}{Z(\beta^{\prime})}=\bigg{\langle}\frac{W(\beta^{\prime\prime})}{W(\beta^{\prime})}\bigg{\rangle}_{MC}$ (1)
where $W(\beta^{\prime})$ and $W(\beta^{\prime\prime})$ denote the weight of the same configuration sampled by MC simulation but with different parameters $\beta^{\prime}$ and $\beta^{\prime\prime}$. In this way, the ratio of partition functions with different parameters can be simulated via Monte Carlo in principle.
@@
Figure 1: Schematic diagram of the reweight-annealing method. (a) Key point 1: Find an annealing path in the whole parameter space to connect the unknown point we want to solved and a known point. $T$ means temperature and $s$ is a general parameter of the Hamiltonian. (b) Key point 2: Use a sampled distribution (red) to represent another distribution (blue) through reweighting a same configuration. If these two distributions are close to each other, the effect of the reweighting would be good since the importance sampling can approximately be kept.
For the first consideration, we take $\beta$ as the inverse temperature. Notice that $Z(0)=d^{N}$ is pre-knowledge because all the configurations are equally important at infinite high temperature, where $N$ and $d$ are the number of particles and the freedom of a single particle respectively. Then we can obtain any $Z(\beta^{\prime})$ from Eq. (1) by setting $\beta^{\prime\prime}=0$. Therefore, the first key point of the method is to choose a known point $Z(\beta^{\prime\prime})$ as the reference which connects to some unknown point $Z(\beta^{\prime})$, as shown in Fig. 1 (a). The parameter of the path is not limited to the temperature $T=1/\beta$, once the parameter path can connect to a known point in the whole parameter space, then the method would work.
The second key point is keeping the importance sampling. Obviously, if the $\beta^{\prime}$ and $\beta^{\prime\prime}$ are close enough, the ratio of PF is close to 1 and is easy for calculation. Otherwise, the ratio would be close to zero or infinite which can not be well-sampled by MC. It can be easily understood in the reweighting frame as shown in Fig. 1 (b), if we want to use a well-known distribution $Z(\beta^{\prime})=\sum W(\beta^{\prime})$ to calculate another distribution $Z(\beta^{\prime\prime})=\sum W(\beta^{\prime\prime})$ via resetting the weight of samplings, the weight before/after resetting of the same configuration should be close to each other. It is still an importance sampling in this sense and requires the $\beta^{\prime}$ and $\beta^{\prime\prime}$ be close enough. As we will discuss later, the practical calculation of Eq. (1) should be divided into many steps for better accuracy, and the inverse temperature $\beta$ will slowly decrease from our concerned point $\beta^{\prime}$ to zero, which has some analogousness with the thermal or quantum annealing algorithm.
When we are considering the vicinity of the zero temperature, the inverse temperature span would be quite wide hence require a long runtime on computer. Thus we can find another alternative parameter path for the calculation according to the first key point mentioned above. One choice in this case is to fix the temperature as a constant. Instead, we introduce an extra term, e.g., transverse field $H_{1}=\sum_{i}\sigma_{i}^{x}$ to the target system $H_{0}$, and consider the partition function $Z^{\prime}(s)$ for $H(s)=sH_{0}+(1-s)H_{1}$, which is exactly the archetypal Hamiltonian of quantum annealing. Similarly, we define ${Z^{\prime}(s^{\prime\prime})}/{Z^{\prime}(s^{\prime})}=\langle{W(s^{\prime\prime})}/{W(s^{\prime})}\rangle_{MC}$ as our observable. The introduction of an external field is conducive because $H(0)$ only has one ground state, thus the entropy $S$ is minimized to be zero and $\ln Z^{\prime}(0)=-\lim_{\beta\to\infty}\beta E$, where $E=-N$ is the ground state energy of $H_{1}$.
Algorithm details.- In this paper, without loss of generality, we focus on the temperature case $\beta=1/T$, the counterpart of the thermal annealing, for demonstrations. We use the stochastic series expansion (SSE) in the followings, in which the weight ratio of temperature is ${W(\beta^{\prime\prime})}/{W(\beta^{\prime})}=({\beta^{\prime\prime}}/{\beta^{\prime}})^{n}$, where $n$ is the the order of series in the expansion. Since the value of the exact zero in the ratio is meaningless, practically, we set $\beta^{\prime\prime}$ in Eq. (1) a sufficiently small number $\beta_{0}$. The adequacy can be checked by comparing the results with that of a smaller $\beta_{0}$ so as to see whether the results ($\ln Z$) converge. Since the estimator in the right hand side of Eq. (1) is unbiased, the algorithm itself is unbiased. Attention that too small ratio of $(\beta_{0}/\beta^{\prime})^{n}$ not only requires more Monte Carlo time $N_{\text{MC}}$ since the error is proportional to $1/\sqrt{N_{\text{MC}}}$, but will exceed the data range of computer (e.g. $1.7\times 10^{-308}\sim 1.7\times 10^{308}$ for double type data).
To fix this problem, according to the second key point mentioned above, we divide $[\beta_{0},\beta^{\prime}]$ into $m$ segments, i.e.
$\frac{Z(\beta_{0})}{Z(\beta^{\prime})}=\prod_{k=1}^{m}\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\prod_{k=1}^{m}\gamma(Z;\beta_{k},\beta_{k-1})$ (2)
where $\beta_{m}\equiv\beta^{\prime}$. We evaluate each $\gamma(Z;\beta_{k},\beta_{k-1})$, and further group them into some batches. It also obviously reveals the method is naturally suitable for parallel computing. The value of each batch is the multiplication of all $\gamma(Z;\beta_{k},\beta_{k-1})$ in it, which we require to be in the same order of magnitude (e.g. $10^{-10}$ in our paper). This can avoid the loss of precision because finally we have to add all the logarithmic values of these batches.
For the multiplication in each batch, to ensure the preciseness, we define the schedule parameter $\alpha_{k}\equiv\beta_{k}/\beta_{k-1}>1$, and require
$\gamma(Z;\beta_{k},\beta_{k-1})=(\frac{1}{\alpha_{k}})^{n}\sim\epsilon$ (3)
According to the computation complexity of SSE a straightforward choice is to set the order of magnitude of $n\sim\mathcal{O}(\beta_{k}L)$ for typical 1D systems and $\mathcal{O}(\beta_{k}L^{2})$ for 2D systems. We can select an empirical and appropriate prefactor $\xi$, setting $n=\xi\beta_{k}L$ for 1D systems, then we do not need to determine $n$ from the thermalization part of Monte Carlo simulation, and the calculations for different $\gamma(Z;\beta_{k},\beta_{k-1})$ can be parallelized. The $\beta_{k}$ here plays similar role as the schedule parameter (temperature or the strength of magnetic field) in annealing algorithms which should vary slow enough to ensure ‘adiabaticity’ and keep the condition of importance sampling. If $\beta_{k}$ varies too dramatically, the accuracy of result will be influenced.
Attention that $n\geq 1$ in Eq. (3), therefore $\alpha_{k,\max}=1/\epsilon$. For low temperature points and large systems, $\alpha_{k}$ can be very close to $1$. To keep more significant digits, a lower bound for $\alpha_{k}$ is also needed, e.g., $\alpha_{k,\min}=1.00001$ in our simulations.
Figure 2: For $\beta=30$, the values of $\ln Z$ as a function of the chain length $L$. Our reweight-annealing (Re-An) method has perfect match with ED. For example, the ED and our results are 251.62283 and 251.624(2) respectively for $L=12$.
1D spin-1/2 Heisenberg chain.- As the first example, we consider a 1D spin-1/2 antiferromagnetic (AFM) Heisenberg chain with periodic boundary condition (PBC) at inverse temperature $\beta=30$. We set the $\epsilon=10^{-2}$ in Eq. 3. For the system size we simulated, the $\beta$ is large enough for checking the effectiveness of our method in low temperature. As it is shown in Fig. 2, results from our method can match that of the exact diagonalization (ED) method exceptionally well for small lengths in a very high-precision. The error bar is about 0.002 while the value is 251.624, and the relative error is smaller than $10^{-7}$. The result is also consistent with the expectation that $\ln Z\propto L+o(L)$, where $o(L)$ means higher order terms which is negligible compared with $L$. One may feel this example too trivial and easy to be calculated. Therefore we introduce the second example which needs a high-precision to extract the correction of the sub-leading term in entropy.
@@
Figure 3: (a) Thermodynamic entropy $S(L)$ as a function of $L$ from different algorithms; (b) The $\ln L$ term extracted from calculating $2S(L)-S(2L)$, and the fitting slope (dashed green line) is 1.00(2).
2D LBW-HM.- To further exploit the advantage of our method, we calculate the thermodynamic entropy of a 2D lattice Bisognano-Wichmann Heisenberg model (LBW-HM) on a square cylinder Giudici et al. (2018); Dalmonte et al. (2018). The thermodynamic entropy $S$ of the LBW-HM at $\beta=1$ provides a good approximation for the Von Neumann entanglement entropy of a 2D PBC square lattice spin-1/2 AFM Heisenberg model under smooth bipartition, i.e., the concerned subsystem is a cylinder Mendes-Santos et al. (2019, 2020).
@@
According to the CFT prediction and previous numerical results Metlitski and Grover (2015); Laflorencie (2016); Zhao et al. (2022a); Deng et al. (2023), the coefficient of sub-leading term of the entropy reflects the number of Goldstone modes. In this example, coarse result of entropy is disfavored because the sub-leading term in entropy is sensitive to errors. For this kind of problems, our method has great advantage over the method of specific heat integral ($C$-int) and Wang-Landau (WL) algorithm Wang and Landau (2001); Troyer et al. (2003) due to its high accuracy and efficiency.
We consider the same LBW-HM as Ref. Mendes-Santos et al. (2020); Giudici et al. (2018) which highly resembles the entanglement Hamiltonian of a spin-1/2 AFM Heisenberg model on a torus with cornerless cutting. Theoretical investigations predict the thermal entropy of the LBW-HM at $\beta=1$ obeys $S(L)=aL+b\ln L+c$, and $b=n_{G}/2$, where $n_{G}=2$ is the number of Goldstone modes for the 2D spin-1/2 AFM Heisenberg model Metlitski and Grover (2011); Laflorencie (2016). Therefore, we can extract $b$ from the slope of equation $2S(L)-S(2L)$, in which the leading term of $L$ can be cancelled with each other.
@@
In Fig. 3, we compare the performance of different algorithms. Generally, the accuracy of the $C$-int method depends on how many intervals we divide. To compare with our method, we require the division of $\beta$ in the $C$-int method to be exactly the same with our method. As it is shown in Fig. 3, this introduces great systemic error because the division number in our method is not adequate for $C$-int. Therefore, to achieve a better result via $C$-int, more computation resources are needed compared to our method.
The WL algorithm, on another hand, suffers from the problem of error saturation, which though can be alleviated by some means like the $t^{-1}$ modification, where $t$ is the Monte Carlo time, the cut-off of the series expansion and the initial value of the refinement parameter $f$ are both hyperparameters to be determined Belardinelli et al. (2008); Belardinelli and Pereyra (2007); Zhou and Bhatt (2005); Inglis and Melko (2013). Therefore, it is very difficult to achieve results with an arbitrary precision with this method and the convergence time is not generally well-controlled. Besides, in Wang-Landau sampling, the change of $f$ typically obeys some monotonically decreasing function, and the computation subroutines with different $f$’s cannot be parallelized because of the linear logic of the algorithm. In this paper, we consider the vanilla WL algorithm, following Ref. Troyer et al. (2003), as an example. Fig. 3(a) shows that the WL algorithm gives roughly correct result for small sizes while the extraction of entropy in Fig. 3(b) completely reveals its unpreciseness, and the $b$ fitted from it is far from the theoretical prediction.
@@
For convenience, here we use equal distance of $T=1/\beta=0.01$ to implement our Re-An method, and it is also easy for numerical differentiation. Because a constant factor of the PF does not affect the (higher order) derivative of $F$, thus we choose $Z(T^{*}=1.95)$ as the reference to replace annealing to a known point. The PF we gained is actually $Z(T)/Z(T^{*})$ where $Z(T^{*})$ is an unknown constant. Then we indeed observe a divergent peak of the second order derivative of the free energy as shown in Fig. 4(a), which accurately probes the phase transition point, matching with the intersection point of binder ratios $R_{2}=\langle m^{4}\rangle/\langle m^{2}\rangle^{2}$ of different sizes shown in Fig. 4(b).
This example also inspires us that connecting to a known point is not necessary when we just need the second or higher order derivative of free energy. The data also demonstrate the high-precision of our method, even the second order derivative is still smooth and precise.
Summary.- We propose an unbiased algorithm within Monte Carlo framework, which has low-technical barrier and no systemic error. The method can be easily implemented for parallel computation and can obtain the value of partition function, free energy and thermal entropy with high precision. The method has two keynotes: 1. Find a parameter path in whole space (either temperature or external parameter) connecting a known point and the to-be-solved one. If we only want to calculate the (high order) derivative of the PF, even fixing an unknown point as the reference is enough; 2. Annealing along the path to calculate the final result through a propagation formula.

### v3

@@ Abstract
On the other hand, the high-precision PF and its related quantities, such as free energy and entropy, are significantly important in numerical calculations since it means more effective information. For phase transitions no matter within or beyond the Ginzburg-Landau paradigm, an efficient and accurate way to straightforwardly obtain PF would be crucial, because its (high order) derivative reveals potential phase transitions. The thermal entropy also relates to the distinctiveness of many novel condensed matters, such as spin ice Snyder et al. (2001); Skjærvø et al. (2020); Udagawa et al. (2021); Bramwell and Harris (2020); Udagawa et al. (2021), topological order Jiang et al. (2012); Castelnovo et al. (2012); Sachdev (2018); Levin and Wen (2006); Yan et al. (2021, 2022), fracton phases Pretko et al. (2020); Nandkishore and Hermele (2019); Gromov and Radzihovsky (2024); Zhou et al. (2022), etc. Besides, the precision of entropy can even reveal important intrinsic physics. In conformal field theory (CFT), the prefactor in front of the sub-leading term of entropy often carries universal information Calabrese and Lefevre (2008); Fradkin and Moore (2006); Nussinov and Ortiz (2009a, b); Casini and Huerta (2007); Ji and Wen (2019, 2020); Kong et al. (2020); Wu et al. (2021a); Zhao et al. (2021); Wu et al. (2021b). Even a small numerical error may make our concerned quantities deviate a lot because the sub-leading term is much more sensitive than the leading one.
@@
In this paper, we present a more accurate and efficient method for calculating PF, dubbed reweight-annealing algorithm, within the framework of Monte Carlo (MC) simulation, either quantum or classical. It is inspired by thermal and quantum annealing algorithm in the quantum simulation area Huse and Fisher (1986); Santoro et al. (2002); Lucas (2014); Kadowaki and Nishimori (1998); Santoro et al. (2002); Mézard and Montanari (2009); Heim et al. (2015); Yan et al. (2023); Ding et al. (2023) and recent developments of the incremental trick in the entanglement calculations of QMC Alba (2017); D’Emidio (2020); Zhao et al. (2022a, b); Pan et al. (2023); Zhou et al. (2024). Our method not only has high accuracy for calculating PF, but is unbiased, easy to implement and naturally can be parallelized on computer.
@@
Basic idea.- Given a partition function $Z({\mathbf{p}})$, where ${\mathbf{p}}=[p_{1},p_{2},\cdots]$ denotes the parameters in the PF such as the temperature $T$ and some Hamiltonian parameter $s$. For example, Fig.1 (a) shows a diagram of the two parameters. For two different parameter points ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}^{\prime\prime}$, we define the reweighting ratio which can be measured in Monte Carlo
$\frac{Z({\mathbf{p}}^{\prime})}{Z({\mathbf{p}}^{\prime\prime})}=\bigg{\langle}\frac{W({\mathbf{p}}^{\prime})}{W({\mathbf{p}}^{\prime\prime})}\bigg{\rangle}_{{\mathbf{p}}^{\prime\prime}}$ (1)
where $W({\mathbf{p}}^{\prime})$ and $W({\mathbf{p}}^{\prime\prime})$ denote the weight of the same configuration under different parameters ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}^{\prime\prime}$, and the subscript indicates the result comes from the MC simulation under parameter ${\mathbf{p}}^{\prime\prime}$. As an example, we take ${\mathbf{p}}=\beta$, where $\beta$ denotes the inverse temperature. Under the framework of stochastic series expansion (SSE), we can then derive $Z(\beta^{\prime})/Z(\beta^{\prime\prime})=\langle(\beta^{\prime}/\beta^{\prime\prime})^{n}\rangle_{\beta^{\prime\prime}}$, where $n$ is the number of non-identity operators in the SSE simulation. The details of the $Z(\beta^{\prime})/Z(\beta^{\prime\prime})$ have been explained in the supplemental materials (SM).
@@
Figure 1: Schematic diagram of the reweight-annealing method. (a) Finding a reference point and the annealing path connecting the unknown point to it in the parameter space. $T$ and $s$ correspond to the thermal and quantum version, respectively; (b) Using a sampled distribution (red) to represent another distribution (blue) through reweighting a same configuration. If these two distributions are close to each other, the effect of the reweighting would be good since the importance sampling can approximately be kept.
The second step is to identify a reference point ${\mathbf{p}}_{0}$, of which the value of $Z({\mathbf{p}}_{0})$ has already been known. Therefore, by evaluating $Z({\mathbf{p}}_{0})/Z({\mathbf{p}}^{\prime\prime})$, we can directly determine the value of $Z({\mathbf{p}}^{\prime\prime})$. In the example that ${\mathbf{p}}=\beta$, apparently $Z(0)=d^{N}$ is pre-knowledge because all the configurations are equally important at infinite high temperature, where $N$ and $d$ are the number of particles and the degree of freedom of a single particle respectively.
However, when ${\mathbf{p}}_{0}$ and ${\mathbf{p}}^{\prime\prime}$ are far away in the parameter space, the value of $Z({\mathbf{p}}_{0})/Z({\mathbf{p}}^{\prime\prime})$ would be close to zero or infinity, resulting in the problem of inefficiency and low precision for the MC simulation. This can be easily understood in the reweighting frame as shown in Fig. 1 (b). If we want to use a well-known distribution $Z({\mathbf{p}}^{\prime\prime})=\sum W({\mathbf{p}}^{\prime\prime})$ to calculate another distribution $Z({\mathbf{p}}^{\prime})=\sum W({\mathbf{p}}^{\prime})$ via resetting the weight of samplings, the weight before/after the resetting of the same configuration should be close to each other. It is still an importance sampling in this sense and requires that ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}^{\prime\prime}$ should be close enough.
Based on this consideration, practically, we divide the calculation of of Eq. (1) into many steps to enhance precision, i.e.
$\frac{Z({\mathbf{p}}_{0})}{Z({\mathbf{p}}^{\prime\prime})}=\prod_{k=1}^{m}\frac{Z({\mathbf{p}}_{k-1})}{Z({\mathbf{p}}_{k})}$ (2)
where ${\mathbf{p}}_{m}\equiv{\mathbf{p}}^{\prime\prime}$ and $Z({\mathbf{p}}_{k-1})/Z({\mathbf{p}_{k}})\to 1$ as long as ${\mathbf{p}}_{k}\to{\mathbf{p}}_{k-1}$. Here, ${\mathbf{p}}$ can be perceived to gradually change from the point of interest ${\mathbf{p}}^{\prime\prime}$ to the reference point ${\mathbf{p}}_{0}$ along a trajectory (annealing path) in the parameter space, akin to an adiabatic process or annealing as shown in Fig.1 (a). Therefore by calculating Eq. (2), not only we are able to achieve a higher precision of $Z({\mathbf{p}}^{\prime\prime})$, but concurrently track the variation of PF along this annealing path. This allows us to further analyze the variation of the free energy (and its derivatives) of the system.
Apparently, the annealing path that ${\mathbf{p}}$ follows determines how large $m$ should we need and how ${\mathbf{p}}$ varies to perform the calculations of Eq. (2). It can be expected that, similar to conventional thermal annealing or quantum annealing, should the annealing trajectory approach a critical point, an increased number of divisions would be necessary in its vicinity. However, we point out that this actually affects little to our algorithm since numerical simulations only deal with finite system, where the criticality does not result in significant inefficiency. We have also verified this point in our simulations. For example, in the 2D spin-$1/2$ XXZ model which we present below, where ${\mathbf{p}}=T$ is the temperature, we implemented an equal division strategy with step size $\Delta T=0.01$. With $20000$ Monte Carlo samples for each bin, we are able to achieve a sufficiently smooth curve for further numerical differentiation. More discussions about the segmentation are in the section of “Dynamical annealing process”.
Thermal and quantum reweight-annealing- As it shown in Fig. 1, if ${\mathbf{p}}=\beta$, as previously mentioned, the annealing path connects an unknown inverse temperature point $\beta^{\prime\prime}$ to the infinite temperature point $\beta_{0}$, which serves as the reference. For convenience, we call this kind of process the thermal reweight-annealing (T-Re-An). Similarly, we can consider the so-called quantum reweight-annealing (Q-Re-An) by setting ${\mathbf{p}}=s$, where $s$ comes from $H(s)=sH_{0}+(1-s)H_{1}$, which is the archetypal Hamiltonian of quantum annealing with $Z(s)=e^{-\beta H(s)}$. Here, $H_{1}$ correspond to our reference point $s=0$, and we would like to obtain the PF related to $H_{0}$. We emphasis that different from the thermal version, the choice of $H_{1}$ here is much more flexible. For example, we can take $H_{1}=\sum_{i}\sigma_{i}^{x}$. Then at the zero temperature ($\beta\to\infty$), $H_{1}$ only has a unique ground state, thus its thermal entropy is minimized to be zero, resulting in $\ln Z(0)=-\lim_{\beta\to\infty}\beta E$, where $E=-N$ is the ground state energy of $H_{1}$. More details about the T-Re-An and Q-Re-An scheme can be found in the SM.
Dynamical annealing process.- For the thermal and quantum reweight-annealing schemes, we provide a dynamical strategy to automatically regulate the alteration of ${\mathbf{p}}$ throughout the annealing process. This not only serves as an economic way for the division in Eq. (2), but can avoid the problem of precision loss when different $r({\mathbf{p}}_{k-1},{\mathbf{p}}_{k})$ differ by several orders of magnitude in the computer. For demonstration, we only discuss the thermal case, i.e. ${\mathbf{p}}=\beta$, in this section.
Without loss of generality, we assume $\beta_{k-1}<\beta_{k}$ in Eq. (2), and define the rate $\alpha_{k}\equiv\beta_{k-1}/\beta_{k}<1$. Within the framework of SSE, we have
$\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\langle\alpha^{n_{k}}_{k}\rangle_{\beta_{k}}$ (3)
where $n_{k}$ is the number of all non-identity operators in the SSE simulation, and the subscript $\beta_{k}$ denotes that this simulation is performed at the parameter point $\beta_{k}$.
The annealing process starts from $k=m$, and we require $Z(\beta_{k-1})/Z(\beta_{k})=\epsilon<1$, where the choice of $\epsilon$ should match $N_{\text{MC}}$, the number of MC steps we use for samplings, as the error in MC simulations is proportional to $1/\sqrt{N_{\text{MC}}}$. Next, we set $n_{k}\approx\tilde{n}_{k}=\Lambda_{k}\beta_{k}L$ ($\Lambda_{k}\beta_{k}L^{2}$) for 1D (2D) system, and $\Lambda_{k}$ can be ascertained through the examination of result convergence. This comes from the fact that in SSE, the energy satisfies $E_{k}=-\langle n_{k}\rangle/\beta_{k}$. Practically, we set $\Lambda_{k}$ to be the same for all $k$ for convenience. This strategy not only makes $Z(\beta_{k-1})/Z(\beta_{k})\approx\epsilon$ for all $k$, but determines the next step in the annealing path, i.e. $\beta_{k-1}=\alpha_{k}\beta_{k}$. The division will be more sparse when $\beta$ gets smaller as we expected, where the thermal fluctuation has destroyed any order and $Z(\beta_{k-1})/Z(\beta_{k})\to 1$, which has little contribution to the result. Apparently, this strategy naturally allows for parallelization of evaluating different $Z(\beta_{k-1})/Z(\beta_{k})$ in Eq. (2). It is also easy to calculate that the number of divisions is around $(\beta-\beta_{0})\Lambda L/|\ln\epsilon|$ for large $L$ or $(\beta-\beta_{0})$, where $\Lambda_{k}=\Lambda$ for all $k$. Therefore, this dynamical annealing strategy does not require exponentially increasing number of divisions.
Figure 2: For $\beta=30$, the values of $\ln Z$ as a function of the chain length $L$, where both T-Re-An method and Q-Re-An method have a perfect match with ED. For example, for $L=12$, the results from the two methods are 251.625(2) and 251.624(2), respectively, and the exact result is 251.62283 from ED calculations. Attention that here in practical SSE simulations, the Hamiltonian has a energy shift of $-1/4$ for each two-body term.
1D spin-1/2 Heisenberg chain.- As the first example, we consider an 1D spin-1/2 antiferromagnetic (AFM) Heisenberg chain with periodic boundary condition (PBC) with $\beta=30$ and length $L$. We use both the T-Re-An method and the Q-Re-An method to calculate the PF. For the quantum version, we write
$H(s)=s\bigg{[}\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}+(1-s)\bigg{[}\sum_{i\in\text{even}}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}$ (4)
where $H_{0}=\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}$ and $H_{1}=\sum_{i\in\text{even}}{\mathbf{S}}_{i}\cdot{\mathbf{S}}_{i+1}$. Apparently, the unique ground state of $H_{1}$ is a tensor product of $L/2$ singlets, each carrying energy $-3/4$ ($\hbar=1$). The estimator for $Z(s^{\prime})/Z(s^{\prime\prime})$ is $\langle(s^{\prime}/s^{\prime\prime})^{n_{\text{odd}}}\rangle_{s^{\prime\prime}}$ in this model, with $n_{\text{odd}}$ denoting the amount of non-identity operators acting on all the bonds with odd indices within the framework of SSE. As it is shown in Fig. 2, both the results of T-Re-An and Q-Re-An method can match that of the exact diagonalization (ED) method exceptionally well for small lengths, which verify the validity of our methods. The results are also consistent with the expectation that $\ln Z\propto L+o(L)$, where $o(L)$ means higher order terms which is negligible compared with $L$. One may feel this example too trivial and easy to be calculated. Therefore we introduce the second example which needs a high-precision to extract the correction of the sub-leading term in entropy.
@@
Figure 3: (a) Thermodynamic entropies $S(L)$ as a function of $L$ calculated from different algorithms; (b) The $\ln L$ term extracted by calculating $2S(L)-S(2L)$, and the fitting slope (dashed green line) here is 1.00(2).
@@
2D LBW-HM.- To further exploit the advantage of our method, we calculate the thermodynamic entropy of a 2D lattice Bisognano-Wichmann Heisenberg model (LBW-HM) Giudici et al. (2018); Dalmonte et al. (2018), the thermodynamic entropy $S$ of the which at $\beta=1$ provides a good approximation for the Von Neumann entanglement entropy of a 2D PBC square lattice spin-1/2 AFM Heisenberg model under smooth bipartition Mendes-Santos et al. (2019, 2020). According to the CFT prediction and previous numerical results Metlitski and Grover (2015); Laflorencie (2016); Zhao et al. (2022a); Deng et al. (2023), the coefficient of sub-leading term of the entropy reflects the number of Goldstone modes. Here, the thermal entropy of the LBW-HM at $\beta=1$ should obey $S(L)=aL+b\ln L+c$, and $b=n_{G}/2$, where $n_{G}=2$ is the number of Goldstone modes of the 2D spin-1/2 AFM Heisenberg model Metlitski and Grover (2011); Laflorencie (2016). In this context, coarse result of entropy is disfavored because the the coefficient $b$ in front of the sub-leading term is sensitive to errors. We have compared our method with the specific heat integral ($C$-int) and Wang-Landau (WL) algorithm in this example.
@@
Similar to our method, the $C$-int method also depends on how many intervals we divide for $\beta$. Therefore we take the same division and Monte Carlo steps to compare the performance of these two algorithms. As it shown in Fig. 3, the $C$-int method has great systemic error because the number of division is adequate for our method, but not the $C$-int method, indicating more computational resources needed for it to achieve better accuracy.
For the WL algorithm, on another hand, the allocation of total computational resources should not be predetermined as it hinges on when we can achieve a converged and flat histogram. Fig. 3 shows great error of the results obtained from the vanilla WL samplings Troyer et al. (2003). This reflects the problem of error saturation, which is intrinsic in this algorithm due to an approximately flat histogram and the empirical refinement parameter Belardinelli et al. (2008); Belardinelli and Pereyra (2007); Zhou and Bhatt (2005). Another disadvantage of the WL sampling compared with our method is that it cannot be parallelized: the two subroutines for flattening the corresponding histograms before and after an adjustment of the refinement parameter, must be executed sequentially. To achieve an arbitrary precision within the Wang-Landau framework is not easy. In contrast, our method is much more accessible as we only need to measure the observable of $Z({\mathbf{p}}_{k-1})/Z({\mathbf{p}}_{k})$ in many conventional and mature MC algorithms.
@@
For convenience, here we use equal distance of $T=1/\beta=0.01$ to implement our Re-An method, and it is also easy for numerical differentiation. Because a constant factor of the PF does not affect the (higher order) derivative of $F$, thus we choose $Z(T^{*}=1.95)$ as the reference to replace annealing to a known point. The PF we gained is actually $Z(T)/Z(T^{*})$ where $Z(T^{*})$ is an unknown constant. Then we indeed observe a divergent peak of the second order derivative of the free energy as shown in Fig. 4(a), which accurately probes the phase transition point, matching with the intersection point of Binder ratios $R_{2}=\langle M^{4}\rangle/\langle M^{2}\rangle^{2}$ of different sizes shown in Fig. 4(b), where $M$ denotes the magnetization.
This example also inspires us that connecting to a known point is not necessary when we just need the second or higher order derivative of free energy. The data also demonstrate the high-precision of our method, where the second order derivative is still smooth and precise.
Summary.- We propose an unbiased algorithm within Monte Carlo framework, which has low-technical barrier and no systemic error. The method can be easily implemented for parallel computation and can obtain the value of partition function, free energy and thermal entropy with high precision. The method has two keynotes: (1) Find a parameter path in whole space (either temperature or external parameter) connecting a known point and the to-be-solved one. If we only want to calculate the (high order) derivative of the PF, even fixing an unknown point as the reference is enough; (2) Annealing along the path to calculate the final result through a propagation formula.